OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 29 vehicles and 82 tires. How many motorcycles passed his house? 6 10 12 17

OpenStudy (johnweldon1993):

Well you know that in total...29 vehicles passed his house...so c + m = 29 We also know that cars have 4 wheels and motorcycles have 2 wheels...and there were a total of 82 tires...so 4c + 2m = 82 Now can you solve the system of equations?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

No don't just guess... c + m =29 4c + 2m = 82 We want to solve for how many motorcycles passed by his house...so lets solve our first equation for 'c' c + m = 29 How would we solve this for 'c'...?

OpenStudy (anonymous):

suub oth sidees

OpenStudy (johnweldon1993):

Subtract what from both sides?

OpenStudy (anonymous):

29

OpenStudy (johnweldon1993):

Not quite...we want to isolate 'c' on the left side....so we would subtract 'm' from both sides...what do we get when we do this?

OpenStudy (johnweldon1993):

c + m = 29 -m -m c = 29 - m right?

OpenStudy (anonymous):

yup

OpenStudy (johnweldon1993):

Now that we know what 'c' equals....we can use it to solve for 'm' in the second equation... 4c + 2m = 82 We know that 'c' = 29 - m...so replace it 4(29 - m) + 2m = 82 now simplify that...can you?

OpenStudy (johnweldon1993):

*hint* distribute the 4 into the parenthesis